OpenStudy (anonymous):

Solve 5x + 2y = 7 10x + 4y = 14 (1, 1) (2, ) No Solution Infinitely Many Solutions

OpenStudy (anonymous):

(2,-3/2) for the second one

OpenStudy (anonymous):

No Solution

OpenStudy (jiteshmeghwal9):

When \[\frac{a_1}{a_2}=\frac{b_1}{b_2}=\frac{c_1}{c_2}\]then there are infinitely solutions